OpenStudy (anonymous):

The graph shows the functions f(x), p(x), and g(x): Graph of function g of x is y is equal to 1.8 to the power of x. The straight line f of x joins ordered pairs 6, 2 and 2, minus 2 and is extended on both sides. The straight line p of x joins the ordered pairs 3, 6 and 6, 2 and is extended on both sides. Part A: What is the solution to the pair of equations represented by p(x) and f(x)? (3 points) Part B: Write any two solutions for f(x). (3 points) Part C: What is the solution to the equation p(x) = g(x)? Justify your answer. (4 points)

OpenStudy (anonymous):

@Compassionate @Hero @iambatman

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

im realllllllllly stuck on this dont even no were to begin and have more if you could help.

OpenStudy (anonymous):

not at all.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

g of x is y is equal to 1.8 to the power of x. \[ g(x) = y = 1.8^x \]

OpenStudy (anonymous):

i need part a for right now.

OpenStudy (anonymous):

what part are you telling me @phi

OpenStudy (phi):

oh, I was starting with the first line. But g(x) is not used in part A.

OpenStudy (anonymous):

im very confused.

OpenStudy (phi):

For f(x), I would write down the two points, and then find the slope between them can you do that ?

OpenStudy (phi):

f of x joins ordered pairs 6, 2 and 2, minus 2 f(x) connects (6,2) and (2,-2) what is the slope ? (i.e. change in y divided by change in x ? )

OpenStudy (anonymous):

can you explain which part your doing like part a = blah blah blah blah that way?

OpenStudy (anonymous):

i dont even know what the points are for f(x)

OpenStudy (phi):

Part A: What is the solution to the pair of equations represented by p(x) and f(x)? to answer, we first need to find the equations. I am starting with finding f(x)

OpenStudy (phi):

they tell you the points, and I typed them in the post up above.

OpenStudy (anonymous):

the points are (6,2)

OpenStudy (anonymous):

hmmmmm....

OpenStudy (phi):

f of x joins ordered pairs 6, 2 and 2, minus 2 f(x) connects (6,2) and (2,-2) what is the slope ? (i.e. change in y divided by change in x ? )

OpenStudy (anonymous):

i dont know how to do slope

OpenStudy (anonymous):

need major help

OpenStudy (phi):

do you know how to find the y value from the (x,y) pair ? what is the y value of (6,2) ?

OpenStudy (anonymous):

i dont really know anything

OpenStudy (anonymous):

its 2

OpenStudy (phi):

and the y value of the other point ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

now find the difference in y (i.e. subtract first minus second)

OpenStudy (anonymous):

4

OpenStudy (phi):

2 - -2 = 2+2 = 4 that is the change in y can you now find the change in x ? same idea, but use the x-values

OpenStudy (anonymous):

so 6-2

OpenStudy (anonymous):

that equals 4

OpenStudy (phi):

yes. now remember slope is "change in y" divided by change in x (say that 3 times to yourself) and then find the slope for these points

OpenStudy (anonymous):

how

OpenStudy (phi):

you found the numbers, now divide them

OpenStudy (anonymous):

this is all i have so far part a= f(x) points are 6,2 and p(x)= 2,-2 then 2-(-2 =4, 6-2=4 part b= part c=

OpenStudy (anonymous):

so i divide 4 and 4 that equals 1

OpenStudy (phi):

yes. and we have not gotten to p(x). so you should write it more as I posted it up above. f(x) is a line with point (6,2) and (2,-2)

OpenStudy (anonymous):

what im so confuse4d.

OpenStudy (anonymous):

please tell me what should be for part a im screwing everything up because thats what i got out of your stuff.

OpenStudy (phi):

They tell us f(x) is a line, that goes through those 2 points.

OpenStudy (anonymous):

what all should be up for part a. like if this was your question that your answering what all would you have right now.

OpenStudy (anonymous):

like right it as part a = blah blah blah please im very lost.

OpenStudy (phi):

How about this. The first is simplify the words. The straight line f of x joins ordered pairs 6, 2 and 2, minus 2 and is extended on both sides. f(x) is a line with points (6,2) and (2,-2) The straight line p of x joins the ordered pairs 3, 6 and 6, 2 and is extended on both sides. write this as p(x) is a line with points (3,6) and (6,2) ok so far ?

OpenStudy (anonymous):

so thats what youd write for part a

OpenStudy (phi):

the solution is the point that is on both lines. In this problem, they tell us a point that is on both lines, so that is the answer.

OpenStudy (anonymous):

part a= f(x)=(6,2) and (2,-2) p(x)= (3,6) and (6,2)

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

but its asking for the solltuions

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Hero